# Saliency strikes back: How filtering out high frequencies improves white-box explanations

### v3

@@ Abstract
Attribution methods correspond to a class of explainability methods (XAI) that aim to assess how individual inputs contribute to a model’s decision-making process. We have identified a significant limitation in one type of attribution methods, known as “white-box” methods. Although highly efficient, these methods rely on a gradient signal that is often contaminated by high-frequency noise. To overcome this limitation, we introduce a new approach called “FORGrad”. This simple method effectively filters out noise artifacts by using optimal cut-off frequencies tailored to the unique characteristics of each model architecture. Our findings show that FORGrad consistently enhances the performance of already existing white-box methods, enabling them to compete effectively with more accurate yet computationally demanding “black-box” methods. We anticipate that our research will foster broader adoption of simpler and more efficient white-box methods for explainability, offering a better balance between faithfulness and computational efficiency.
@@ 1 Introduction
Understanding how modern Artificial Neural Networks (ANNs) make decisions remains a major challenge. Given the ever-increasing range of machine learning applications, the need for robust and reliable explainability methods is pressing (Doshi-Velez & Kim, 2017a; Jacovi et al., 2021). To meet this demand, eXplainable Artificial Intelligence (XAI) focuses on developing new tools to help users better understand how ANNs arrive at their decision (Jacovi et al., 2021; Doshi-Velez & Kim, 2017b; Rudin, 2019). Among the wide range of XAI approaches, attribution methods have become the go-to approaches. They aim to identify the features that drive a network’s decisions (Simonyan et al., 2013) by assigning them an importance score according to their contribution to the overall prediction.
Attribution methods fall broadly into two main classes: “white-box” and “black-box” methods. White-box methods require full access to a network’s information (weights, architecture, etc.) to understand which regions in the input are the most important for the model’s prediction. Typically, such methods rely on the gradient of the model’s prediction w.r.t. the input to produce attribution maps. For instance, “Saliency”, one of the earliest white-box methods, traces a model’s decision back to its input image using such gradient information (Simonyan et al., 2013). Although simple and computationally efficient, Saliency is also known to produce noisy and hard-to-interpret attribution maps (Nguyen et al., 2021; Kim et al., 2022; Colin et al., 2021; Serrurier et al., 2022; Hase & Bansal, 2020). Several extensions have been proposed to try to overcome these limitations, e.g., by varying slightly the input and accumulate the resulting gradient to obtain a smoother map (Simonyan et al., 2013; Zeiler & Fergus, 2014; Springenberg et al., 2014; Smilkov et al., 2017; Sundararajan et al., 2017; Zhou et al., 2016; Selvaraju et al., 2017a; Fong & Vedaldi, 2017).
Conversely, black-box methods focus on a model’s prediction without explicitly requiring any gradient computation. Such methods aim to quantify how changes in the inputs affect a model’s output by altering the input and evaluating the newly obtained prediction (Zeiler & Fergus, 2014; Petsiuk et al., 2018; Fel et al., 2021a; Ribeiro et al., 2016; Novello et al., 2022). These black-box methods are computationally much more demanding as they require hundreds of network queries (or forward passes), but they outperform white-box methods on current XAI benchmarks. We refer the reader to Section 2 for more details on the different types of attribution methods. Despite extensive research (Sixt et al., 2020; Adebayo et al., 2018; Slack et al., 2021; Han et al., 2022; Agarwal et al., 2022, 2021; Pukdee et al., 2023), a question remains: Why do black-box methods outperform white-box methods?
@@
Consistent with findings from multiple studies (Kim et al., 2019; Springenberg et al., 2014; Selvaraju et al., 2017b; Rao et al., 2022; Adebayo et al., 2018; Tomsett et al., 2019; Ghorbani et al., 2017; Ancona et al., 2018), we have found empirically that white-box methods typically produce attribution maps with increased power in the high frequencies relative to their black-box counterparts. Our analysis further reveals that these high frequencies are inherited from the gradient itself. We have also found that these high frequencies often arise due to some downsampling operations within the network, such as max-pooling or striding. This observation suggests that white-box methods could be “repaired” by filtering out the gradient signal with carefully chosen network-specific cut-off frequencies.
Here, we introduce FORGrad, FOurier Reparation of the Gradients, a simple method applicable to all white-box methods with negligible computational overhead. FORGrad filters high-frequencies in the gradient using a cut-off frequency adapted to the network architecture and the attribution method. This frequency is precomputed on training images to maximize the faithfulness of individual attribution methods. Experimentally, we demonstrate that FORGrad consistently improves all existing white-box attribution methods. The method and its effects are summarized in Fig. 1.
@@
  • We found that a major source of high-frequency noise in attribution maps computed with white-box methods is inherited from the model gradient. This noise is responsible for their lower explainability score.
  • These gradient artifacts are a consequence of the max-pooling or striding operations used in convolutional neural networks (CNNs), and removing them does not affect the model prediction.
  • We introduce FORGrad to repair white-box methods. FORGrad filters-out frequencies above a certain cut-off value which is optimized for a given network architecture. FORGrad systematically improves the explainability score of white-box methods to levels where they now compete with the computationally intensive black-box methods.
@@ 2 Related Work
Gray-box attribution methods: The distinction between white-box and black-box methods is to be nuanced, particularly as some approaches leverage only partial knowledge of the network. Grad-CAM (Selvaraju et al., 2017a) and Grad-CAM++ (Chattopadhay et al., 2018) fall within this category. These methods use the gradient information only of the last layers of the network, as opposed to the full back-propagation done by white-box methods. Specifically, Grad-CAM and Grad-CAM++ generate visual explanations for class labels using a weighted combination of the positive partial derivatives of the last convolutional layer feature maps with respect to a specific class score. This process generates small sample-specific maps that are then upsampled to fit the input size. Although these methods perform relatively well and are computationally efficient, the corresponding attribution maps often lack fine details (Nam et al., 2021). This limitation primarily stems from the low-spatial resolution gradient maps they rely on to produce attribution maps.
@@ 3 Notations, Metrics, and Networks
Notations: We consider a general supervised learning setting, where a classifier $\bm{f}:{\mathcal{X}}\to{\mathcal{Y}}$ maps images from an input space ${\mathcal{X}}\subseteq\mathbb{R}^{W\times H}$ to an output space ${\mathcal{Y}}\subseteq\mathbb{R}$. We denote $\mathcal{F}$ the Discrete Fourier Transform (DFT) on $\mathbb{R}^{W\times H}$. Importantly, all Fourier spectra shown in this article are symmetrized, i.e. we always shift the low-frequency components to the center of the spectrum. We recall that an attribution method is a function $\bm{\varphi}:{\mathcal{X}}\to\mathbb{R}^{W\times H}$ that maps an input of interest to its corresponding importance scores $\bm{\varphi}({\bm{x}})$. Finally, we denote by $\bm{\varphi}_{\sigma}({\bm{x}})$ the attribution method in which high frequencies are filtered out using a cut-off value of $\sigma$.
@@
XAI metrics: To evaluate XAI performance, we use the faithfulness metric, denoted $F(\bm{x},\bm{\varphi})$, which is a measure that includes the Insertion and Deletion metrics (Petsiuk et al., 2018). Those metrics are widely used in XAI to evaluate the quality of attribution methods (Fong & Vedaldi, 2017). The Deletion metric, $D(\bm{x},\bm{\varphi})$, measures the evolution of the prediction probability when one incrementally removes pixels by replacing them with a baseline value according to their attribution score. Insertion, $I(\bm{x},\bm{\varphi})$, inserts pixels gradually into a baseline input. The faithfulness is obtained using $F(\bm{x},\bm{\varphi})=I(\bm{x},\bm{\varphi})-D(\bm{x},\bm{\varphi})$. Note that the faithfulness metric has shown to be effective in evaluating attribution methods (Li et al., 2022). Additionally, we also evaluate attribution methods using the $\mu$Fidelity, denoted $\mu$(Bhatt et al., 2020). The fidelity metric serves to verify the correlation between the attribution score and a random subset of pixels. To achieve this, a set of pixels is randomly chosen and set to a baseline state, after which a prediction score is obtained. The fidelity metric evaluates the correlation between the decrease in the score and the significance of the explanation for each random subset created.
@@ 4 White-box methods are contaminated by high-frequencies from the gradient
Figure 3: High-frequency power in attribution methods. A: Fourier signature of white-box and black-box attribution methods. White-box methods produce attribution maps with increased power in high frequencies. B: Faithfulness and power/frequency slope for several attribution methods. The numbers below the bar correspond to this method’s computational time, measured using a set of 100 images (ImageNet) with an Nvidia T4. White-box methods show lower faithfulness but significantly better computational complexity compared to black-box methods. See section 3 for more details on how faithfulness, the power/frequency slope, and the Fourier signature were computed.
@@ 4.1 Comparison between white-box and black-box methods
White-box methods are computationally more efficient but have lower faithfulness: We quantify the computational time of each method to generate an explanation on 100 ImageNet images with an Nvidia T4. The computational time is reported for each attribution method using numbers placed under their corresponding bar in Fig. 3B. We refer the reader to section 3 for more details on the faithfulness computation. Overall, Fig. 3B demonstrates that white-box (and gray-box) methods are computationally more efficient than black-box methods. In addition, it shows that white-box methods have a lower power/frequency slope and a lower faithfulness than black-box methods. This suggests that the presence of high-frequency noise in white-box methods might be a limiting factor in achieving XAI performance similar to black-box methods.
@@ 4.2 A low-pass filtered gradient approximates well the original gradient
White-box methods produce attribution maps by relying on the gradients of the classifier w.r.t. the input, denoted $\nabla_{\bm{x}}\bm{f}(\bm{x})$. The prevalence of high frequencies in these attribution maps suggests that the gradients harbor high-frequency noise. To study such a noise in the gradients, we leverage the first-order approximation of the model, i.e., $\bm{f}(\bm{x}+\bm{\bm{\varepsilon}})\approx\bm{f}(\bm{x})+\bm{\varepsilon}\nabla_{\bm{x}}\bm{f}(\bm{x})$, and we compute the $\ell_{2}$ approximation error, over different values of $\bm{\varepsilon}$.
@@
Figure 4: Gradients are contaminated with high frequencies and stem from MaxPooling operations. A: We evaluate the importance of high-frequency content in the gradient using a first-order approximation of the model, i.e. $\bm{f}(\bm{x}+\bm{\varepsilon})\approx\bm{f}(\bm{x})+\bm{\varepsilon}\nabla_{\bm{x}}\bm{f}(\bm{x})$, and we compute the $\ell_{2}$ approximation error when we remove high frequency up to $\sigma$: $\zeta({\bm{x}},\sigma)$ relative to the error when we do not filter $\zeta({\bm{x}},\sigma_{\text{max}})$. Gradients in which we remove high-frequency content (dark blue) produce an error closer to the baseline (green), compared to the control conditions (pink). B: Example of gradients in a ResNet architecture after a pooling operation, along with their Fourier signature. MaxPooling operations elicit high-frequency power in the gradient.
4.3 High-frequency artifacts stems from max-pooling operation
@@
Our investigation now turns to understanding why high-frequency artifacts appear in gradients. In Fig. 4B, we analyze the power spectrum of various gradients in a ResNet50, along with their visualization. Consistent with findings from Olah et al. (2017), we observe a clear checkerboard pattern in the gradients following a max-pooling. This pattern is reduced after an average-pooling due to a different mechanism in the backward implementation of both operations (see Fig. A.3 for more visualizations). Their respective Fourier signature (Fig. 4B right) confirm our hypothesis: the max-pooling operation shows more power in the high frequencies compared to the average-pooling (see Fig. A.4 for all pooling layers). We additionally investigate the contribution of every layer in the frequency power of the gradients. We analyze the power/frequency slope of gradients across various layers of ResNet50 (see Fig. A.5). We observe a jump in the power/frequency slope at each max-pooling operation, in between ResNet blocks (see blue curve in Fig. A.5). To discern whether this increase in high frequency is a result of the network architecture or the training process, we conduct a similar analysis with an untrained ResNet50 (black curve in Fig.A.5) and observe a similar trend. Subsequently, we replace max-pooling with average-pooling to ensure the preservation of information continuity in the gradient (red curve in Fig.A.5). We observe a clear divergence in the power/frequency slope between max-pooling and average-pooling networks, highlighting the crucial role of max-pooling in creating high-frequency signals in gradients. We observe a similar pattern of results with a VGG16 (Simonyan & Zisserman, 2014) architecture (see Figures A.6,A.7, A.8).
@@
In summary, we have established that the high-frequency noise contaminating white-box methods originate from the gradient. Specifically, max-pooling operations within the network are the primary source of this high-frequency noise. These insights underscore the potential benefits of implementing a low-pass filter, with a cut-off frequency adapted to the network architecture to repair white-box attribution methods.

### v4

@@ Abstract
Attribution methods correspond to a class of explainability methods (XAI) that aim to assess how individual inputs contribute to a model’s decision-making process. We have identified a significant limitation in one type of attribution methods, known as “white-box” methods. Although highly efficient, as we will show, these methods rely on a gradient signal that is often contaminated by high-frequency artifacts. To overcome this limitation, we introduce a new approach called “FORGrad.” This simple method effectively filters out these high-frequency artifacts using optimal cut-off frequencies tailored to the unique characteristics of each model architecture. Our findings show that FORGrad consistently enhances the performance of existing white-box methods, enabling them to compete effectively with more accurate yet computationally more demanding “black-box” methods. We anticipate that, because of its effectiveness, the proposed method will foster the broader adoption of straightforward and efficient white-box methods for explainability, providing a better balance between faithfulness and computational efficiency.
@@ 1 Introduction
Understanding how modern Artificial Neural Networks (ANNs) make decisions remains a major challenge. Given the ever-increasing range of machine learning applications, the need for robust and reliable explainability methods is pressing (Doshi-Velez & Kim, 2017a; Jacovi et al., 2021). To meet this demand, eXplainable Artificial Intelligence (XAI) focuses on developing new tools to help users better understand how ANNs arrive at their decisions (Jacovi et al., 2021; Doshi-Velez & Kim, 2017b; Rudin, 2019). Among the wide range of XAI approaches, attribution methods have become the go-to approaches. They aim to identify the features that drive a network’s decisions (Simonyan et al., 2013) by assigning them an importance score according to their contribution to the overall prediction. Attribution methods fall broadly into two main classes: “white-box” and “black-box” methods. White-box methods require full access to a network’s information (weights, architecture, etc.) to understand which regions in the input are the most important for the model’s decision. Typically, such methods rely on the gradient of the model’s prediction w.r.t. the input to produce attribution maps. For instance, “Saliency”, one of the earliest white-box methods, traces a model’s decision back to its input image using such the gradient information (Simonyan et al., 2013). Although simple and computationally efficient, Saliency is also known to produce noisy and hard-to-interpret attribution maps (Nguyen et al., 2021; Kim et al., 2022; Colin et al., 2021; Serrurier et al., 2022; Hase & Bansal, 2020). Several extensions have been proposed to try to overcome these limitations, e.g., by varying slightly the input and accumulating the resulting gradient to obtain a smoother map (Simonyan et al., 2013; Zeiler & Fergus, 2014; Springenberg et al., 2014; Smilkov et al., 2017; Sundararajan et al., 2017; Zhou et al., 2016; Selvaraju et al., 2017a; Fong & Vedaldi, 2017).
Conversely, black-box methods focus on a model’s predictions without explicitly requiring any gradient computation. Such methods aim to quantify how changes in the inputs affect a model’s output by altering the input and evaluating the newly obtained prediction (Zeiler & Fergus, 2014; Petsiuk et al., 2018; Fel et al., 2021a; Ribeiro et al., 2016; Novello et al., 2022). These black-box methods are computationally much more demanding as they require hundreds of network queries (or forward passes), but they outperform white-box methods on current XAI benchmarks. We refer the reader to Section 2 for more details on the different types of attribution methods. Despite extensive research (Sixt et al., 2020; Adebayo et al., 2018; Slack et al., 2021; Han et al., 2022; Agarwal et al., 2022, 2021; Pukdee et al., 2023), a question remains: Why do black-box methods outperform white-box methods?
@@
Consistent with findings from multiple studies (Springenberg et al., 2014; Selvaraju et al., 2017b; Ghorbani et al., 2017; Ancona et al., 2018; Balduzzi et al., 2017; Adebayo et al., 2018; Kim et al., 2019; Tomsett et al., 2019; Rao et al., 2022), we have found empirically that white-box methods typically produce attribution maps with increased power in the high frequencies relative to their black-box counterparts. Our analysis further reveals that these high frequencies are inherited from the gradient itself. We have also found that these high frequencies often arise due to some downsampling operations within the network, such as max-pooling or striding. This observation suggests that white-box methods could be “repaired” by filtering out the gradient signal with carefully chosen network-specific cut-off frequencies.
Here, we introduce FORGrad  FOurier Reparation of the Gradients, a simple method applicable to all white-box methods with negligible computational overhead. FORGrad filters high-frequencies in the gradient using a cut-off frequency adapted for the network architecture and the attribution method used. This frequency is precomputed on training images to maximize the faithfulness of individual attribution methods. Experimentally, we demonstrate that FORGrad consistently improves all existing white-box attribution methods. The method and its effects are summarized in Fig. 1.
@@
  • We found that a major source of high-frequency artifacts in attribution maps computed with white-box methods is inherited from the model gradient. These artifacts are responsible for the lower explainability scores of these methods.
  • These gradient artifacts are a consequence of the max-pooling and striding operations used in convolutional neural networks (CNNs).
  • We introduce FORGrad to repair white-box methods. FORGrad filters-out frequencies above a certain cut-off value, which is optimized for a given network architecture. FORGrad systematically improves the explainability score of white-box methods to levels where they now compete with the computationally intensive black-box methods.
@@ 2 Related Work
Gray-box attribution methods: The distinction between white-box and black-box methods is nuanced, particularly as some approaches leverage only partial knowledge of the network. Grad-CAM (Selvaraju et al., 2017a) and Grad-CAM++ (Chattopadhay et al., 2018) fall within this category. These methods use the gradient information only of the last layers of the network, as opposed to the full back-propagation done by white-box methods. Specifically, Grad-CAM and Grad-CAM++ generate visual explanations for class labels using a weighted combination of the positive partial derivatives of the last convolutional layer feature maps with respect to a specific class score. This process generates small sample-specific maps that are then upsampled to fit the input size. Although these methods perform relatively well and are computationally efficient, the corresponding attribution maps often lack fine details (Nam et al., 2021). This limitation primarily stems from the low-spatial resolution gradient maps they rely on to produce attribution maps.
@@ 3 Notations, Metrics, and Networks
Notations: We consider a general supervised learning setting, where a classifier $\bm{f}:{\mathcal{X}}\to{\mathcal{Y}}$ maps images from an input space ${\mathcal{X}}\subseteq\mathbb{R}^{W\times H}$ to an output space ${\mathcal{Y}}\subseteq\mathbb{R}$. We denote $\mathcal{F}$ the Discrete Fourier Transform (DFT) on $\mathbb{R}^{W\times H}$. Importantly, all Fourier spectra shown in this article are symmetrized, i.e., we always shift the low-frequency components to the center of the spectrum. We recall that an attribution method is a function $\bm{\varphi}:{\mathcal{X}}\to\mathbb{R}^{W\times H}$ that maps an input of interest to its corresponding importance scores $\bm{\varphi}({\bm{x}})$. Finally, we denote by $\bm{\varphi}_{\sigma}({\bm{x}})$ the attribution method in which high frequencies are filtered out using a cut-off value of $\sigma$.
@@
XAI metrics: To evaluate XAI performance, we use the faithfulness metric, denoted $F(\bm{x},\bm{\varphi})$, which is a measure that includes the Insertion and Deletion metrics (Petsiuk et al., 2018). Those metrics are widely used in XAI to evaluate the quality of attribution methods (Fong & Vedaldi, 2017). The Deletion metric, $D(\bm{x},\bm{\varphi})$, measures the evolution of the prediction probability when one incrementally removes pixels by replacing them with a baseline value according to their attribution score. Insertion, $I(\bm{x},\bm{\varphi})$, inserts pixels gradually into a baseline input. The faithfulness is obtained using $F(\bm{x},\bm{\varphi})=I(\bm{x},\bm{\varphi})-D(\bm{x},\bm{\varphi})$. Note that the faithfulness metric has shown to be effective in evaluating attribution methods (Samek et al., 2016; Li et al., 2022). Additionally, we also evaluate attribution methods using the $\mu$Fidelity, denoted $\mu$(Bhatt et al., 2020). The fidelity metric serves to verify the correlation between the attribution score and a random subset of pixels. To achieve this, a set of pixels is randomly chosen and set to a baseline state, after which a prediction score is obtained. The fidelity metric evaluates the correlation between the decrease in the score and the significance of the explanation for each random subset created. Finally, the sensitivity metric, $S(\bm{x},\bm{\varphi})$ (Bhatt et al., 2020), measures a method’s stability by calculating the distance between two explanations. These explanations are generated by the same method, with the first using the original input and the second using the same input perturbed with random noise. A good method, according to this metric, should produce similar explanations for the two inputs, therefore obtaining a low sensitivity score.
@@ 4 White-box Methods are Contaminated by High-Frequencies from the Gradient
Figure 3: High-frequency power in attribution methods. A: Fourier signature of white-box and black-box attribution methods. White-box methods produce attribution maps with increased power in the high frequencies. B: Faithfulness and power/frequency slope for several attribution methods. The numbers below the bar correspond to this method’s computational time, measured using a set of 100 images (ImageNet) with an Nvidia T4. White-box methods show lower faithfulness but significantly better computational complexity compared to black-box methods. See section 3 for more details on how faithfulness, the power/frequency slope, and the Fourier signature were computed.
@@ 4.1 Comparison between White-Box and Black-Box Methods
White-box methods are computationally more efficient but have lower faithfulness: We quantify the computational time of each method to generate an explanation on 100 ImageNet images with an Nvidia T4. The computational time is reported for each attribution method using numbers placed under their corresponding bar in Fig. 3B. We refer the reader to section 3 for more details on the faithfulness computation. Overall, Fig. 3B demonstrates that white-box (and gray-box) methods are computationally more efficient than black-box methods. In addition, it shows that white-box methods have a lower power/frequency slope and a lower faithfulness than black-box methods. This suggests that the presence of high-frequency artifacts in white-box methods might be a limiting factor in achieving XAI performance similar to black-box methods.
@@ 4.2 A Low-Pass Filtered Gradient Approximates Well the Original Gradient
White-box methods produce attribution maps by relying on the gradients of the classifier w.r.t. the input, denoted $\nabla_{\bm{x}}\bm{f}(\bm{x})$. The prevalence of high frequencies in these attribution maps suggests that the gradients harbor high-frequency artifacts. To study such artifacts in the gradients, we leverage the first-order approximation of the model, i.e., $\bm{f}(\bm{x}+\bm{\bm{\varepsilon}})\approx\bm{f}(\bm{x})+\bm{\varepsilon}\nabla_{\bm{x}}\bm{f}(\bm{x})$, and we compute the $\ell_{2}$ approximation error, over different values of $\bm{\varepsilon}$.
@@
Figure 4: Gradients are contaminated with high frequencies and stem from MaxPooling operations. A: We evaluate the importance of high-frequency content in the gradient using a first-order approximation of the model, i.e., $\bm{f}(\bm{x}+\bm{\varepsilon})\approx\bm{f}(\bm{x})+\bm{\varepsilon}\nabla_{\bm{x}}\bm{f}(\bm{x})$, and we compute the $\ell_{2}$ approximation error when we remove high frequency up to $\sigma$: $\zeta({\bm{x}},\sigma)$ relative to the error when we do not filter $\zeta({\bm{x}},\sigma_{\text{max}})$. Gradients in which we remove high-frequency content (dark blue) produce an error closer to the baseline (green) compared to the control conditions (pink). B: Example of gradients in a ResNet architecture after a pooling operation, along with their Fourier signature. MaxPooling operations elicit high-frequency power in the gradient.
4.3 High-Frequency Artifacts Stems from Max-Pooling Operations
@@
Our investigation now turns to understanding why high-frequency artifacts appear in gradients. In Fig. 4B, we analyze the power spectrum of various gradients in a ResNet50, along with their visualization. Consistent with findings from Olah et al. (2017), we observe a clear checkerboard pattern in the gradients following a max-pooling. This pattern is reduced after an average-pooling due to a different mechanism in the backward implementation of both operations (see Fig. A.3 for more visualizations). Their respective Fourier signature (Fig. 4B right) confirm our hypothesis: the max-pooling operation shows more power in the high frequencies compared to the average-pooling (see Fig. A.4 for all pooling layers). We additionally investigate the contribution of every layer in the frequency power of the gradients. We analyze the power/frequency slope of gradients across various layers of ResNet50 (see Fig. A.5). We observe a jump in the power/frequency slope at each max-pooling operation in between ResNet blocks (see blue curve in Fig. A.5). To discern whether this increase in high frequency is a result of the network architecture or the training process, we conduct a similar analysis with an untrained ResNet50 (black curve in Fig.A.5) and observe a similar trend. Subsequently, we replace max-pooling with average-pooling to ensure the preservation of information continuity in the gradient (red curve in Fig.A.5). We observe a clear divergence in the power/frequency slope between max-pooling and average-pooling networks, highlighting the crucial role of max-pooling in creating high-frequency signals in gradients. We observe a similar pattern of results with a VGG16 (Simonyan & Zisserman, 2014) architecture (see Figures A.6,A.7, A.8). Additionally, as noted by Odena et al. (2016), this effect can be mitigated by adjusting the stride value during downsampling. Consequently, in Fig. A.9, we examine the impact of striding on high-frequency power in a VGG16 network.
@@
In summary, we have established that the high-frequency artifacts contaminating white-box methods originate from the gradient. Specifically, max-pooling operations within the network appear to be a primary source of these high-frequency artifacts. These insights underscore the potential benefits of implementing a low-pass filter with a cut-off frequency adapted to the network architecture to repair white-box attribution methods.